OpenStudy (twizttiez):

Janine inflated 1 ball to a radius of 18 cm and another ball to a radius of 12 cm. How much more air was in the larger ball? Use 3.14 to approximate pi and express your answer in hundredths. ??? cm^3

OpenStudy (twizttiez):

@Mehek14

Mehek (mehek14):

\(V-v = \frac{ 4 }{ 3 }\pi r^{3} - \frac{ 4 }{3 } \pi r^{3}\)

OpenStudy (twizttiez):

Okaii one second ^-^

Mehek (mehek14):

find the volume for both and subtract

OpenStudy (twizttiez):

I got 24429.02447431 for the first one

Mehek (mehek14):

ok good now for the second one

OpenStudy (twizttiez):

Wait would the exact form be 7776 pi?

OpenStudy (twizttiez):

And for 12 I got 7238.22947387

Mehek (mehek14):

lets see B1 = 24429.02 B2 = 7238.23 24429.02 - 7238.23 =

OpenStudy (twizttiez):

Oh okaiii one more step and I'll be done! XD

OpenStudy (twizttiez):

17100.79?

Mehek (mehek14):

yup

OpenStudy (twizttiez):

Thanks!